PLEASE TYPE THE ANSWER

(14%) Y = -2X2

  1. (10%) Compute / derive the PDF f Y(y) for Y = -2X2 when X is uniformly distributed over [-1, 3]. Show the range where f Y(y) is nonzero, i.e. [a, b], what is a and what is b?

(b) 4%) Verify the PDF f Y(y) you calculated in (a) is a PDF
Answer #1

We have the PDF of X is fx (x) = 3; -1 <x<3 .

a) Consider the tranformation Y = -282.

The PDF of the tranformation Y = g(x) is

fy(Y) = { fx (g-?(y)) dg (V)

The summation is over the number of inverse functions.

Here i F = (1), 6 = x . The PDF of Y = -282 is

fr(Y) = 4,-20;-18 <y5-2

Hence [a, b] = [-18, -2]

b) To verify that fY(Y) = 4,27; -18<y<-2 is a PDF,

Lacondy Landovanos -18 1-2 fy(y)dy = 1 J-18

Hence fY(Y) = 4,27; -18<y<-2 is a valid PDF.